5. Consider the initial value problem ?′′+9?′+18?=3? ,?(0)=2,?′(0)=−1
a. Determine an appropriate ?? for the differential equation.
b. Using concepts from “Undetermined Coefficients,” determine the appropriate ??.
c. Using concepts from “Variation of Parameters,” determine the appropriate ??.
d. Using the results from parts a-c, determine the solution to the initial value problem

Please use the specified method, no matrices
Answer #1

☆ Given y+9y+ 18y = 3t 9(0): 2, y(o)=-1 09+90 +18) y = 30 Auxiliary ean is flm) 20 m² tam+ 18 20 m + 6 m+ 3m +18=0 m(m +6)

undertermined Coefficients CO? + +40 +18) y - 3+ 9p= Az + B&te yp = 2AX+B 9. -ZA 2A +9 (24*+B) +18 (4&+B4+c) = 36 Compare t C

initial Condetion : 9(0)=2, y(OJ = -1 = t -6+ -3t y=cé t qe alt 6 12 Ylo) y=2 when t=0, - 2 co ) - + Çeo) 2 the -ام a tę - t

25 12 C+C2 = 띔 L 2C, +% 13 C = 잃 t. 이 36 25 ㅋ 1 -C- 뜷 12 CS SL 61 36 - - 5 alt 25 (2) 156 정 2 3 Ca SL-nl 9. 19 ~ 36 16t -3t e